Thursday, 30 October 2025

Wednesday October 29th

Mike Wise and Ed Ford. Our mission was to clear the debris from last week's Hookah Pipe (HP) diameter adjustment and if needed repeat the process. We declined the suggestion to take in more cement so just took a 36v drill, a couple of 10mm drill bits, a spare set of batteries and passage widening kit in two bags.

A slight whiff of fumes was detected from the pit above the Light Connection (LC). MW descended to check things out while EF waited above. It seemed OK at first; there was a pile of small debris in the HP but further slight adjustment was needed and so the drill bag came down. There was no detectable draught and TPC was the same temperature as the upper cave. But then the fumes became more apparent in the area around HP and upstairs and the drill bag returned up LC. EF came down for a look and chat.  

EF declined to enter HP; his sense of smell hasn't been wrecked by time like mine has. We did discuss moving some rocks towards the platform below LC keeping us away from the lower part of TPC, but then EF traversing the rock pile found a pocket of fume. Time was called and we returned up LC. Despite using exemplary thrutching technique we were both panting on reaching the top and according to EF, MW was sweating buckets.

We exited puffing and panting with throats complaining we had just eaten a handful of gravel. No photos taken. Medicinal substances were taken in The Hunters' which helped to improve things slightly.

1 hour 40 minutes

 Mike Wise

Monday, 27 October 2025

Monday 27th October 2025

Mike Kushy and Tuska. 2 hours fitting metal gate. Need an extension bar for the driver to get the lower screws in on the hinges and need to sort out a lock. Mike.

Gate fitted. Photo MK.

Gate fitted. Photo MK.

Sunday, 26 October 2025

Sunday 26 October 2025

Mike Kushy and Tuska. MK picked up a piece of 3mm sheet metal that John Gisborne had cut for us, thanks JG! The sheet was then taken to Tuska's house where we cut to hole size and attached hinges and handles. Gate now ready to fit! 2 hours. Mike.

Top side of Gate. Photo MK.

Underside of Gate. Photo MK.

Wednesday, 22 October 2025

Wednesday 22nd October 2025

Mike Kushy, Mike Wise and Ed Ford. We met up at 4pm and headed in together. MW and EF went in the usual route carrying bags whilst MK wound up and collected the top wire.,Tthis was taken further down to leap frog the lower wire. MW set to drilling and removed the 1m drill bit so MK could go ahead and grab the newer rope (the air felt quite thick in the intermediate chamber and drill dust was just hovering). Leaving MW to drill shot holes MK and EF replaced the rope on the light connection and set about building the landing area properly with rocks and mortar. 3 lots of sand left in cave and half a cement, loads of water. MW then used what battery was remaining to drill a bolt hole to create a traverse type line (this is only half drilled). MK meanwhile set the charge. We had a little look down 'Ed's bet' (the vertical pot directly below Cement Dig) and it does look worth going in and clearing the rocks out at some point soon. We fired from the top of the Lower Bedding, heard and felt faintly. 3 hours. Mike K.

MW and EF building in TPC. Photo MK.

MW and EF building in TPC. Photo MK.

Sunday, 19 October 2025

Saturday 18th October 2025

Mike Kushy, Hollie London, Mikey Waterworth and Olivia Dawson. MK and HL to begin with started at 10.30am, we cleared the bang spoil from the up and over boulder to start. The space over the top is a lot more comfortable, although the lower piece is also shattered and could be removed entirely (personally I think we should just cement in and stabilise what's there). Next job was to drill the stemples in (thanks to Lee V for use of your Hilti drill for this), one on the side wall below the ledge and another up near the scaff' bar. We could do with putting a couple of smaller metal holds higher up again, but all a lot safer now. Next MK started drilling for shot holes in The Hookah Tube, 3 holes @600 and 1 at just shy of 1m (1m drill bit currently sticking out of the wall, battery ran out!), whilst HL went on into MHH and started to dig out the final squeeze some more. A retaining wall was built at the 'up cave' end of MHH to put all the glorious mud behind. CO2 and O2 tests were done throughout, see results below. Around 12.30 MW and OD came and met us in MHH. They said the air looked very thick. We agreed a plan and continued the already started dig to the right of the formation. Unfortunately this hit a solid wall 1m in and won't give us the clearance to dig around the formation so we will have to try a dig in the far corner? (tarpaulin sheet now over the formation for protection). MK and HL left the others to continue digging as they both felt puffy and like they needed better air. Really felt the air quality change once back in TPC. HL started to rebuild the landing area below the Light Connection. We had a tidy round and got packed to leave around 3pm. MK went through to MHH to let the others know our plan and we all decided to all leave together (also getting close to 3% CO2, our set safety limit). Very wet in MHH today with puddles everywhere. Would recommend not going into MHH for a while and let the air settle, so we can retest the differences (probably the worst day of the year today for testing air quality in a cave, so hopefully it passes). Oh MK and HL went to mix some muck towards the end but realised there was no cement in the cave. About 5 drums of sand though, 10l of mixing water and a bag of stone dust. 5 hours. Mike K. 

MK drilling holes for stemples. Photo HL.

CO2 and O2 results (thanks to Stuart 'Mac' for use of your monitors):

Start (11.30am) - CO2: 2.08, O2: 18.76%

12.01pm - CO2: 2.31, O2: 18.5%

End (14:16) - CO2: 2.56, O2: 18.18%

*Apart from 15 minutes (around 1pm) there was only 2 people in MHH at any time and working whilst we monitored*